OpenStudy (anonymous):

HELP NOW OR A$AP PLEASE I'LL AWARD MEDAL.....What is the length of side BC???

OpenStudy (anonymous):

HERE'S THE PICTURE

OpenStudy (anonymous):

@tejasvir

OpenStudy (anonymous):

or you can use sin30 = x/ 30

OpenStudy (anonymous):

JUST GIVE ME THE CORRECT ANSWER ROUNDED TO ONE DECIMAL PLACE

OpenStudy (anonymous):

wait i mean tan of 30

OpenStudy (ibbutibbu.):

you are not allowed to give answers @gswag98

OpenStudy (anonymous):

yeah i been helping people since I got me account on here

OpenStudy (anonymous):

and I've been getting medals

OpenStudy (ibbutibbu.):

But it is against the code of conduct to give answers unless they are clueless

OpenStudy (anonymous):

I am clueless and trying to hurry up and get the answer typed on my program thingie

OpenStudy (ibbutibbu.):

Im not starting an argument, I am just telling you.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

solve for x do tan30 = x/ 30 so x = 30 tan 30

OpenStudy (anonymous):

ok so what I got was 17.32050808 now what does that look like rounded to one decimal place @tejasvir @Ibbutibbu.

OpenStudy (ibbutibbu.):

17.3

OpenStudy (anonymous):

THANK YOU BOTH!!!! My Program Thingie said "Way To Go"

OpenStudy (anonymous):

yea np:)